OpenStudy (anonymous):

Find the domain and range of x^2-6x+14-y=0

OpenStudy (phi):

I would write this as y= x^2-6x+14 the domain are the "legal" x values if you don't divide by 0 or take the root of a negative number, all x values are legal so you should know what the domain is, right?

OpenStudy (anonymous):

I guess just all real numbers

OpenStudy (phi):

yes (-inf, +inf)

OpenStudy (phi):

for the range, you have to find the min and max of the y values. Notice that this is a parabola. I would find the vertex, and then the min y value. the max y value is ??

OpenStudy (anonymous):

the max y value? infinity right?

OpenStudy (phi):

yes y goes to +inf for polynomials you can just focus on the highest degree x^2 here and ask what happens when x is +(very big number) or -(very big number). you get a +(very big number) which we call +infinity so the range is [ ?, +inf)

OpenStudy (phi):

for ax^2 + bx +c the vertex is at x=-b/(2a) use that x to find the min y value

OpenStudy (anonymous):

the vertex is (3, 5), so y is greater than or equal to 5. i was absent during this lesson, and the class apparently completed the square, to get (x-3)^2=y-5, and then got the answer. how do you get the vertex from there?

OpenStudy (phi):

the equation of a parabola in vertex form is y = a (x-h)^2 +k the vertex is (h,k) (hence the name, "vertex form" in this case y = (x-3)^2 + 5 and you read off (3,5)

OpenStudy (phi):

you can see when x is 3 you get (3-3)^2 or 0 and 0+5 is 5 if x is something other than 3 you get a number and when you square it, it is positive so you add something to 5, and y is bigger. This is a long way to say x=3 must give the min value of y

OpenStudy (phi):

your answer is domain (-inf, +inf) range [5, +inf) (you use square bracket to show the range includes 5) you use parens for the infinity (because infinity isn't a number)

OpenStudy (anonymous):

thanks!